# DTN:  Deep Multiple Task-specific Feature Interactions Network for Multi-Task Recommendation

### v3

@@ 3. $\bullet$
In this section, we conduct extensive offline and online experiments on both our company’s large-scale recommender system and public benchmark datasets. These experiments aim to evaluate the effectiveness of the proposed models. The results demonstrate that our approach outperforms state-of-the-art MTL methods.
Evaluation on the E-commerce Recommender System in Our Company
@@
In this subsection, we perform offline and online experiments on various tasks within our company’s E-commerce recommender system to assess the performance of the proposed models.
Dataset
@@
We collected traffic logs from the online E-commerce recommender system in our company, of which one week’s samples are used for training and samples of the following day for testing. The size of training and testing set is about 6.3 billions and 0.9 billion respectively. We focus on three tasks in this dataset, namely CTR, ATC, and CVR, which model user preferences.
Models Census Task1 Census Task2 AE CTR AE CVR AliCCP CTR AliCCP CVR
AUC RelaImp AUC RelaImp AUC RelaImp AUC RelaImp AUC RelaImp AUC RelaImp
ShareBot 0.9361 - 0.9915 - 0.7173 - 0.8111 - 0.6061 - 0.6121 -
@@
Table 3: Experiment Results on Public Datasets
Models for Comparisons
In the experiment, we compare SFM, TFI and DTN model with baseline PLE model. Considering that DHEN is a serial method which will greatly increase model latency, DTN does not use this feature interaction method. In our DTN model we employed the TSN for the ATC target as well for better performance. We introduce three newest feature interaction methods used in DTN as follows:
  1. $\bullet$
    GDCN: GDCN (Wang et al. 2023) utilizes low-rank methods to approximate feature interactions in subspaces, achieving a better trade-off between model performance and latency. We choose GDCN as a feature interaction module instead of DCN (Wang et al. 2017) and DCN-M(Wang et al. 2020).
  2. $\bullet$
    MemoNet: MemoNet (Zhang and Zhang 2023) use multi-hash codebooks to memorize and learn the feature interaction. In order to keep the same params as other feture interaction methods, we only choose part of user features and item features with top feature importances to cross in MemoNet, which may not result in the optimal performance for this feature interaction method.
  3. $\bullet$
    MaskNet: MaskNet (Wang, She, and Zhang 2021) has been proposed to learn specific representation vectors for each feature considering its context, which we think should be very useful in ranking models in RS.
Offline Experiment Setup
In our experiment, CTR, ATC and CVR prediction task are binary classification tasks, trained by cross-entropy loss and evaluation metric by AUC. Each FIM outputs dimensions in DTN is 512, and task-sensitive network. The output dimension of each feature interaction module in SFM, TFI, and DTN remains consistent at 512. We use Adam(Kingma and Ba 2014) optimizer in our experiment, and the default learning rate is set to 1e-3. The batch size is set to 2048. The embedding dimentions is 16. Unless otherwise specified, all activation functions are ReLU. We use 2-layer gated cross layers for our GDCN interaction. And use part of user and item features for our MemoNet method, with 2 HCNets, and we apply 2-order interactions inside the MemoNet. For MaskNet method, the mask gate block we use a hidden layer dimensions of 640, apply layer normalization to it, and then compress it to the same dimension of 512 at output. We use 12 TFI for each model and control the parameter size of each TFI to be similar. The overall parameter is around 1.6E8 for different models to make a fair comparison, except for $DTN\_trim$ model, which reduced about 9% of parameters than DTN.
Evaluation Metrics
In our experiments, we use AUC (Area Under ROC) and LogLoss as the evaluation metric. AUC has an upper bound of 1, and a higher value indicates better performance. RelaImp, as described in work (Xie et al. 2021), is used as another evaluation metric to measure the relative AUC improvements over the corresponding baseline model. Since AUC is 0.5 from a random strategy, we can remove the constant part of the AUC score and formalize the RelaImp as:
$RelaImp=(\frac{AUC(Measured\ Model)-0.5}{AUC(Base\ Model)-0.5}-1)\times 100\%,$ (8)
Evaluation on our company dataset
@@
The experiment results from our company’s offline evaluations are summarized in Table 2. From Table 2 we can conclude that applying feature interactions such as GDCN, MemoNet and MaskNet in SFM has demonstrated an uplift in AUC metrics. Further enhancements were achieved by applying the task-specific feature interaction module in TFI model, which allowed us to learn the task-specific feature interaction representations. The most significant improvements were observed with our DTN model, in which we integrate three feature interaction methods in this experiment, resulting in the highest metric uplifts for CTR, ATC, and CVR prediction tasks. And we also include a DTN-trim version for comparison, where we utilize gating weights to trim the DTN. This shows the effectiveness of MFI module in DTN, where diversified task-specific feature interaction modules are selectively utilized to optimize performance across varied tasks.
Online A/B Testing
@@
The DTN model has been applied within our company’s E-commerce recommender system and rigorously tested to evaluate its effectiveness in a live environment. We conducted a 7-days online A/B testing to directly compare the performance of DTN against the existing PLE model. The results of this testing are summarized in Table 4, which details the improvements achieved by our model. The results indicate a significant uplift in clicks by +3.28%, and orders + 3.10% and GMV (Gross Merchandise Volume) +2.70%. The result shows the improvements of DTN model compared to state-of-the-art MTL models.
@@
Table 4: Improvement over PLE on Online A/B Test.
Evaluation on Public Datasets
@@ 1. $\bullet$
    Census-Income Dataset(mis 2000) is a public dataset extracted from the 1994 census brureau database, which provide 40 features for training and two tasks for prediction, one task is whether the income exceeds 50K, the other task is to predict whether the person is married.
@@ 2. $\bullet$
    Public AE Dataset(pengcheng Li et al. 2020) is a public dataset which contains 22,326,719 training samples and 9,342,708 test samples gathered from real-world traffic logs of AliExpress. We will predict the CTR and CVR task inside this dataset.
@@ 3. $\bullet$
    Ali-CCP Dataset(Ma et al. 2018a) is a public dataset from Taobao’s traffic logs with 42,300,135 training samples and 43,016,840 test samples included. It also focuses on CTR and CVR predictions.
Implementation Details
The setup for census-income dataset is the same as (Ma et al. 2018a). For AE and Ali-CCP dataset, we use a PLE for each task as base Model to compare with, which is same as (Tang et al. 2020). Same as implement details introduced in Offline Experiment Setup section, we implement all the models with Tensorflow in our experiments and control the overall parameter size of each model to have a fair comparison.
Experiment Results
@@
Combining previous experiments on the industrial dataset and online A/B tests conducted within our company, the DTN model has demonstrated extraordinary performance in learning representations through multiple diversified feature interaction methods, resulting in the most significant performance improvement.
Figure 4: Visualization of FIM Weights for CTR, ATC, and CVR Tasks in DTN.
Feature Interaction Selection
@@
As illustrated in 4, we compare the contributions of MaskNet, MemoNet and GDCN in DTN that trained in our company data. We can observe that for each tasks, their respective MaskNet feature interaction modules contribute the most. And for the ATC task, MemoNet demonstrates the best performance in its own MFI module as well as in the CVR MFI module. For the CVR task, GDCN outperforms MemoNet in its own MFI module.
@@
Figure 5: Visualization of DTN: (a) DTN can learn task-specific representations with same FIM method. (b) Different FIM learn diversified representatioins for same task.
Visualization of DTN
Finally, we visualize the task-specific feature interaction representations of the feature interaction modules in DTN on our company dataset. It demonstrates that DTN can learn diversified task-specific feature interaction representations, which enhances the model’s performance.
We start by visualizing the task-specific feature interaction representations from the feature interaction modules across the CTR, ATC, and CVR tasks on our company dataset. We randomly select 1,000 samples from the dataset, and taking the MaskNet feature interaction module as example, which have a better performance than MemoNet and GDCN as we showed in Table 2. Fig. 5(a) illustrates the visualization of the representation of the MaskNet feature interaction method using t-SNE(Van der Maaten and Hinton 2008) learned by DTN, in which points with same color correspond to the same task. As expected, Fig. 5(a) demonstrates that the representations vectors of task-specific feature interactions exhibit significant differences in comparison to both other tasks and the shared feature interactions. This emphasizes the importance of incorporating task-specific feature interactions.

### v4

@@ 5. Experiment
In this section, we conduct extensive offline and online experiments on both Shopee’s large-scale recommender system and public benchmark datasets. These experiments aim to evaluate the effectiveness of the proposed models. The results demonstrate that our approach outperforms state-of-the-art MTL methods.
5.1. Evaluation on the E-commerce Recommender System in Shopee
@@
In this subsection, we perform offline and online experiments on various tasks within Shopee’s E-commerce recommender system to assess the performance of the proposed models.
5.1.1. Dataset
@@
We collected traffic logs from the online E-commerce recommender system in Shopee, of which one week’s samples are used for training and samples of the following day for testing. The size of training and testing set is about 6.3 billions and 0.9 billion respectively. We focus on three tasks in this dataset, namely CTR, ATC, and CVR, which model user preferences.
5.1.2. Models for Comparisons
In the experiment, we compare SFM, TFI and DTN model with the Shared-Bottom, MMoE and PLE model. Considering that DHEN is a serial method which will greatly increase model latency, DTN does not use this feature interaction method. In our DTN model we employed the TSN for the ATC target as well for better performance. We introduce three newest feature interaction methods used in DTN as follows:
  1. $\bullet$
    GDCN: GDCN (Wang et al., 2023) utilizes low-rank methods to approximate feature interactions in subspaces, achieving a better trade-off between model performance and latency. We choose GDCN as a feature interaction module instead of DCN (Wang et al., 2017) and DCN-M(Wang et al., 2020).
  2. $\bullet$
    MemoNet: MemoNet (Zhang and Zhang, 2023) use multi-hash codebooks to memorize and learn the feature interaction. In order to keep the same params as other feture interaction methods, we only choose part of user features and item features with top feature importances to cross in MemoNet, which may not result in the optimal performance for this feature interaction method.
  3. $\bullet$
    MaskNet: MaskNet (Wang et al., 2021) has been proposed to learn specific representation vectors for each feature considering its context, which we think should be very useful in ranking models in RS.
5.1.3. Offline Experiment Setup
In our experiment, CTR, ATC and CVR prediction task are binary classification tasks, trained by cross-entropy loss and evaluation metric by AUC. Each FIM outputs dimensions in DTN is 512, and task-sensitive network. The output dimension of each feature interaction module in SFM, TFI, and DTN remains consistent at 512. We use Adam(Kingma and Ba, 2014) optimizer in our experiment, and the default learning rate is set to 1e-3. The batch size is set to 2048. The embedding dimentions is 16. Unless otherwise specified, all activation functions are ReLU. We use 2-layer gated cross layers for our GDCN interaction. And use part of user and item features for our MemoNet method, with 2 HCNets, and we apply 2-order interactions inside the MemoNet. For MaskNet method, the mask gate block we use a hidden layer dimensions of 640, apply layer normalization to it, and then compress it to the same dimension of 512 at output. We use 12 TFI for each model and control the parameter size of each TFI to be similar. The overall parameter is around 1.6E8 for different models to make a fair comparison, except for $DTN\_trim$ model, which reduced about 9% of parameters than DTN.
5.1.4. Evaluation Metrics
In our experiments, we use AUC (Area Under ROC) and LogLoss as the evaluation metric. AUC has an upper bound of 1, and a higher value indicates better performance. RelaImp, as described in work (Xie et al., 2021), is used as another evaluation metric to measure the relative AUC improvements over the corresponding baseline model. Since AUC is 0.5 from a random strategy, we can remove the constant part of the AUC score and formalize the RelaImp as:
(8) $RelaImp=(\frac{AUC(Measured\ Model)-0.5}{AUC(Base\ Model)-0.5}-1)\times 100\%,$
Table 3. Experiment Results on Public Datasets
Models Census Task1 Census Task2 AE CTR AE CVR Ali-CCP CTR Ali-CCP CVR
AUC RelaImpr AUC RelaImpr AUC RelaImpr AUC RelaImpr AUC RelaImpr AUC RelaImpr
Shared-Bottom 0.9361 - 0.9915 - 0.7173 - 0.8111 - 0.6061 - 0.6121 -
@@ 5.1.4. Evaluation Metrics
5.1.5. Evaluation on Shopee dataset
@@
The experiment results from Shopee’s offline evaluations are summarized in Table 2. From Table 2 we can conclude that applying feature interactions such as GDCN, MemoNet and MaskNet in SFM has demonstrated an uplift in AUC metrics. Further enhancements were achieved by applying the task-specific feature interaction module in TFI model, which allowed us to learn the task-specific feature interaction representations. The most significant improvements were observed with our DTN model, in which we integrate three feature interaction methods in this experiment, resulting in the highest metric uplifts for CTR, ATC, and CVR prediction tasks. And we also include a DTN-trim version for comparison, where we utilize gating weights to trim the DTN. This shows the effectiveness of MFI module in DTN, where diversified task-specific feature interaction modules are selectively utilized to optimize performance across varied tasks.
5.1.6. Online A/B Testing
@@
The DTN model has been applied within Shopee’s E-commerce recommender system and rigorously tested to evaluate its effectiveness in a live environment. We conducted a 7-days online A/B testing to directly compare the performance of DTN against the existing online model, which is PLE. The results of this testing are summarized in Table 4, which details the improvements achieved by our model. The results indicate a significant uplift in clicks by +3.28%, and orders + 3.10% and GMV (Gross Merchandise Volume) +2.70%. The result shows the improvements of DTN model compared to state-of-the-art MTL models.
Table 4. Improvement over PLE on Online A/B Test.
@@
5.2. Evaluation on Public Datasets
@@ 1. $\bullet$
    Census-Income Dataset(mis, 2000) is a public dataset extracted from the 1994 census brureau database, which provide 40 features for training and two tasks for prediction, one task is whether the income exceeds 50K, the other task is to predict whether the person is married.
@@ 2. $\bullet$
    Public AE Dataset(pengcheng Li et al., 2020) is a public dataset which contains 22,326,719 training samples and 9,342,708 test samples gathered from real-world traffic logs of AliExpress. We will predict the CTR and CVR task inside this dataset.
@@ 3. $\bullet$
    Ali-CCP Dataset(Ma et al., 2018b) is a public dataset from Taobao’s traffic logs with 42,300,135 training samples and 43,016,840 test samples included. It also focuses on CTR and CVR predictions.
5.2.2. Implementation Details
The setup for census-income dataset is the same as (Ma et al., 2018b). For AE and Ali-CCP dataset, we also compare DTN with Shared-Bottom, MMoE and PLE models, which is same as (Tang et al., 2020). Same as implement details introduced in Offline Experiment Setup section, we implement all the models with Tensorflow in our experiments and control the overall parameter size of each model to have a fair comparison.
5.2.3. Experiment Results
@@
Combining previous experiments on the industrial dataset and online A/B tests conducted within Shopee, the DTN model has demonstrated extraordinary performance in learning representations through multiple diversified feature interaction methods, resulting in the most significant performance improvement.
5.3. Feature Interaction Selection
@@
As illustrated in 4, we compare the contributions of MaskNet, MemoNet and GDCN in DTN that trained in Shopee data. We can observe that for each tasks, their respective MaskNet feature interaction modules contribute the most. And for the ATC task, MemoNet demonstrates the best performance in its own MFI module as well as in the CVR MFI module. For the CVR task, GDCN outperforms MemoNet in its own MFI module.
@@
Figure 4. Visualization of FIM Weights for CTR, ATC, and CVR Tasks in DTN.
Figure 5. Visualization of DTN: (a) DTN can learn task-specific representations with same FIM method. (b) Different FIM learn diversified representatioins for same task.
5.4. Visualization of DTN
Finally, we visualize the task-specific feature interaction representations of the feature interaction modules in DTN on Shopee dataset. It demonstrates that DTN can learn diversified task-specific feature interaction representations, which enhances the model’s performance.
We start by visualizing the task-specific feature interaction representations from the feature interaction modules across the CTR, ATC, and CVR tasks on Shopee dataset. We randomly select 1,000 samples from the dataset, and taking the MaskNet feature interaction module as example, which have a better performance than MemoNet and GDCN as we showed in Table 2. Fig. 5(a) illustrates the visualization of the representation of the MaskNet feature interaction method using t-SNE(Van der Maaten and Hinton, 2008) learned by DTN, in which points with same color correspond to the same task. As expected, Fig. 5(a) demonstrates that the representations vectors of task-specific feature interactions exhibit significant differences in comparison to both other tasks and the shared feature interactions. This emphasizes the importance of incorporating task-specific feature interactions.